# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v1

@@ Abstract
The Internet of Medical Things (IoMT) transcends traditional medical boundaries, enabling a transition from reactive treatment to proactive prevention. This innovative method revolutionizes healthcare by facilitating early disease detection and tailored care, particularly in chronic disease management, where IoMT automates treatments based on real-time health data collection. Nonetheless, its benefits are countered by significant security challenges that endanger the lives of its users due to the sensitivity and value of the processed data, thereby attracting malicious interests. Moreover, the utilization of wireless communication for data transmission exposes medical data to interception and tampering by cybercriminals. Additionally, anomalies may arise due to human errors, network interference, or hardware malfunctions. In this context, anomaly detection based on Machine Learning (ML) is an interesting solution, but it comes up against obstacles in terms of explicability and protection of privacy. To address these challenges, a new framework for Intrusion Detection Systems (IDS) is introduced, leveraging Artificial Neural Networks (ANN) for intrusion detection while utilizing Federated Learning (FL) for privacy preservation. Additionally, eXplainable Artificial Intelligence (XAI) methods are incorporated to enhance model explanation and interpretation. The efficacy of the proposed framework is evaluated and compared with centralized approaches using multiple datasets containing network and medical data, simulating various attack types impacting the confidentiality, integrity, and availability of medical and physiological data. The results obtained offer compelling evidence that the FL method performs comparably to the centralized method, demonstrating high performance. Additionally, it affords the dual advantage of safeguarding privacy and providing model explanation.
@@
Internet of Medical Things Intrusion Detection System Machine Learning Federated Learning explainable AI Security Privacy.
@@ 1 INTRODUCTION
Internet of Things is a technology revolutionizing the field of information science by incorporating sensors associated with objects to collect data. Its ability to obtain information anywhere and anytime has led to its integration into various sectors, including the healthcare domain known as IoMT. Equipped with sensors and actuators, medical devices facilitate continuous, remote, and real-time collection of physiological data, such as glucose levels, body temperature, and heart rate, enabling constant health monitoring. The augmentation in both quantity and quality of amassed data serves to optimize treatment efficacy, mitigate medical inaccuracies, and expedite early disease detection. This transformation means a transition from curative to preventive healthcare, which considerably increases patients’ chances of recovery.
@@
Moreover, leveraging the gathered health data, automatic treatments can be administered to patients with chronic conditions throw actuators. For example, diabetic patients can receive automated insulin injections based on their blood glucose levels. Similarly, individuals with irregular heart rhythms can be administered electrical shocks through pacemakers, while those with neurological disorders can benefit from simulated brain activity via Deep Brain Implants, enhancing the overall well-being and quality of life for patients with chronic illnesses. The collected health information can be stored on cloud servers or hospital databases for in-depth analysis, harnessing the power of AI-assisted healthcare under the supervision of healthcare professionals, often referred to as Healthcare 4.0.
@@
Notwithstanding the myriad benefits offered by IoMT, it faces significant security and privacy challenges, as evidenced by the alarming statistic indicating a 77% rise in malware attacks on IoT devices in the first six months of 2022 [1]. The utilization of wireless communication for transmitting data exposes it to man-in-the-middle attacks. Furthermore, other sources of anomalies may affect the integrity of data caused by human errors during data processing, network interference during transmission, or malfunctions occurring at the medical equipment level [2]. These anomalies can compromise the integrity, confidentiality, and availability of crucial medical data. Such anomalies could lead to misdiagnoses and errors in treatment, potentially resulting in tragic consequences.
@@
To mitigate these concerns regarding anomalies, ML has been proposed as a solution. Considering that IoMT systems generate large amounts of data, they can help ML models to distinguish between normal and abnormal behaviour. This capability facilitates the detection of anomalies and is effective against zero-day and new attacks. These detection systems can operate in real-time, and with advancements in Deep Learning (DL), the process of attribute selection contributing to detection becomes automated.
Despite the potential benefits, ML-based security solutions face several challenges. A significant limitation is the integration of IDS-based ML into a centralized architecture, which raises concerns about data privacy. Sharing sensitive information in this manner may violate user privacy, and the central node’s compromise poses a single point of failure risk for the entire system. Moreover, the centralized architecture is susceptible to latency issues as all data must be transmitted to the central node. This limitation can hinder network expansion and potentially lead to a lack of computational capacity for training the model. Additionally, ML encounters the challenge of explaining its outcomes. The reliance on ”black-box” models, where the decision-making process is opaque, makes it difficult to analyze. This issue is particularly relevant in the healthcare domain, where international organizations such as Health Insurance Portability and Accountability Act (HIPAA) mandate demands for explainability and transparency. Meeting these requirements is crucial for stakeholders to comprehend and interpret model predictions, thereby reinforcing trust in the obtained results.
@@
This article aims to introduce a framework to enhance the security of IoMT systems through the design of an IDS based on ML. The proposed solution involves the utilization of FL as a training methodology, allowing the sharing of locally trained model weights on end-devices instead of raw data. This preserves data privacy in alignment with regulations, and the distributed nature of FL mitigates the single point of failure associated with a centralized structure. By opting for the sharing of model weights rather than raw data, numerous key challenges are addressed. This not only diminishes bandwidth consumption but also alleviates network congestion, thereby facilitating system scalability. Adopting FL enhances data confidentiality, reduces the potential risks associated with a centralized model, and increases overall system efficiency and robustness. Furthermore, exploration is undertaken into the utilization of XAI for model explanation and interpretability. This approach seeks to enhance the comprehensibility of AI models for all stakeholders, including patients, model designers, and regulators. Through transparent explanations and a proven track record of reliable performance, this framework cultivates trust among users, facilitating confidence in the predictive capabilities of the system.
The proposed solution’s contributions can be summarized as follows:
@@
    The integration of XAI methods enables the explanation and interpretation of the proposed ML models for both patients and model designers. This not only enhances transparency but also ensures compliance with regulatory requirements and healthcare legislation.
@@
    The proposed solution’s performance was thoroughly evaluated and compared with the centralized method, demonstrating its effectiveness.
@@
    The evaluation of the solution proposed has been conducted on four distinct datasets containing network and medical data, further validating its applicability and robustness.
The article is structured as follows: Section 2 provides background and related work. Section 3 outlines the methodology for proposed IDS model. In Section 4, the experimental setup and results are described. Section 5 includes a discussion of the contribution. Finally, Section 6 draws conclusions from the method and presents future perspectives for the work.
@@ 2.1 Background
Shifting to ML, which can be categorized into supervised, unsupervised, and semi-supervised learning paradigms. Supervised ML involves training models on labeled data for classification or regression tasks, capturing relationships between inputs and output data [4]. Unsupervised learning explores patterns in unlabeled data [5], while semi-supervised learning leverages both labeled and unlabeled datasets to improve model performance and generalization [6]. The response time of an IDS spans the spectrum from real-time detection to a predefined time interval. Upon the detection of an attack, the system initiates a response, which may be passive, involving the notification of the administrator regarding the presence of an attack, or active, entailing actions like communication blockage. The architecture can take different forms, such as Centralized, Fully-Distributed, and Partially Distributed IDS with hierarchical reporting mechanisms.
@@
Introducing FL, a distributed ML approach preserving privacy by transmitting models instead of rows data. In FL, a global model is constructed by leveraging the participation of multiple devices over communication rounds. A subset of clients is chosen at the beginning of each round, updating the model independently using locally stored data. The server aggregates these models to create an enhanced global model, iterating until convergence. Two types of FL emerge: Cross-Silo FL, involving distributed data centers, and Cross-Device FL, encompassing mobile devices and IoT equipment [7]. FL encompasses various categorizations depending on the data distribution characteristics. These include Horizontal FL, where data across multiple devices share identical features but possess distinct samples; Vertical FL, where data across diverse entities exhibit differing features but may share common identifiers; and Federated Transfer Learning, which extends conventional Transfer Learning to the federated setting, allowing the transfer of knowledge across distributed data sources. [8], each contributing to data privacy and learning process improvement.
@@
However, ML models are often considered black-box. To address this, Van Lent introduced XAI in 2004 [9]. XAI comprises methodologies aiding researchers in comprehending and gaining trust in ML model outcomes. A comprehensive AI model can be characterized with a three-dimensional description [10]. Explainability refers to the capacity to articulate the learning model’s processes, interpretability provides insight into the model’s operation, and transparency denotes inherent understandability without user intervention. In the domain of explainability, transparent models, like Decision Tree (DT), embody ante-hoc explainability, while opaque models, like Deep Neural Networks (DNN), require post-hoc explainability. Nevertheless, there are scenarios where even transparent models with seemingly straightforward rules require post-hoc elucidation. Despite the complexity of opaque models, they often outperform transparent ones, leading to a trade-off between performance and explainability. Post-hoc methods operate on both local and global scop, employing techniques such as feature relevance analysis, which assesses the significance of each input feature in the model’s predictions. Surrogate models contribute to interpretability by simplifying the primary model through methods like probing with local changes or utilizing its structure. Additionally, representative examples, drawn from the model’s training set, showcasing significant levels of confidence in their classification towards a particular class [11].
@@ 2.2 Related Work
Initially, centralized IDS solutions utilizing ML for anomaly detection are reviewed. Subsequently, an examination is conducted on solutions employing FL for intrusion detection. Finally, systems of detection based on ML, incorporating XAI, are explored.
@@
The research presented by [14] present a methodology aimed at detecting anomalies and cyberattacks within healthcare systems. Their approach involves optimal feature selection using Recursive Feature Elimination (RFE), a process that iteratively eliminates the least relevant features using Logistic Regression (LR) or XGB Regressor methods. Subsequently, they employ Multi-layer Perceptron (MLP) with parameter optimization for attack detection. Before applying their method, they conducted data preprocessing, replacing NaN values with column means and utilizing label encoding to transform categorical values into numerical ones. Performance evaluation was conducted using various metrics including accuracy, precision, recall, and f1-score. Their methods were applied to IoMT datasets such as WUSTL-EHMS [15], ECU-IoHT [16], Intensive care unit (ICU) [17], and ToN-IoT [18]. Furthermore, they compared their results with previous studies. The findings demonstrate that the combination of XGB Regressor for feature selection and optimized parameter settings for MLP yielded the most promising results among the tested methodologies.
In the work by [19] have introduced an IDS designed specifically for connected healthcare systems. Their proposed methodology involves the utilization of Stacked Sparse Autoencoder (sSAE) to reduce dimensionality and the memory required for computing the covariance matrix. Subsequently, they employ the Sliced Gated Recurrent Unit (SGRU), a parallelized version of the RNN, achieved through processing segmentation. To measure the performance of their solution, the authors employed the AWID dataset [20]. However, recognizing the imbalance in the dataset, they solved this problem by applying the Synthetic Minority Oversampling Technique (SMOTE) to augment the original dataset with synthetic data. In their comparative analysis, the authors benchmarked their solution against Deep Neural Network (DNN), Random Forest (RF), Long Short Term Memory (LSTM), LA-SMOTE-GRU [21], Auto Encoder (AE), XGBoost hybrid model, and LSTM hybrid model. Notably, their results demonstrate that their proposed solution outperforms these existing models, achieving superior performance metrics while simultaneously reducing model size and processing time.
@@
The authors of [27] have introduced an IDS tailored for the IoMT. Their innovative methodology incorporates PSO for feature selection followed by Deep Neural Network (DNN) utilization for intrusion detection within their proposed system. In their evaluation, the authors tested their solution on a dataset encompassing both network and medical features. To comprehensively assess their approach, they compared the results of their PSO-DNN solution with various ML and DL models, including LR, KNN, DT, RF, SVM, CNN, and LSTM on WUSTL-EHMS datasets [15]. For this comparative analysis, accuracy, precision, recall, and f1-score were utilized as evaluation metrics. Their findings revealed that the DL models surpassed the performance of traditional ML models. Furthermore, their PSO-DNN solution not only outperformed the ML models but also surpassed the state-of-the-art approaches in intrusion detection within the IoMT domain.
@@
The IDS proposed in [31] is tailored for detecting data flow modifications within multi-cloud healthcare systems, comprising gateways, edge cloud, and a core cloud. Their proposed approach involves leveraging Deep Hierarchical Stacked Neural Networks. This methodology entails reusing layers trained at the edge cloud level and merging them at the core cloud level to create a pre-trained model. To validate their solution, they conducted testing on UNSW-BOT-IoT [32] and UNSW-NB15 datasets [26], including one generated by the authors. A comparative analysis was performed between their model, which incorporates reusing the trained layers, and a model that does not. Their findings indicate that their solution enhances accuracy while reducing training time.
@@
Finally, the comparative study by [33] emphasizes the impact of combining network and medical attributes for intrusion detection in IoMT. Initially, they undertook the creation and collection of both medical and network data, along with simulating a man-in-the-middle attack. Subsequently, they tested various ML algorithms, including ANN, RF, KNN, and SVM, based on metrics such as AUC, accuracy, and execution time. The results indicate that the performance of ML algorithms utilizing the combination of network and medical attributes yields superior results compared to using them separately.
@@
Continuing our exploration, attention now turns towards IDS solutions harnessing the power of FL. A noteworthy example is found in the work of [34], where the authors present a framework for storing and privacy protection of data used to detect cyber-attacks within IoMT systems. Their approach involves distributing heterogeneous data across distributed cloud nodes with privacy protection measures. They further suggested merging data from diverse sources using the differential privacy contractive deep autoencoder. This technique facilitates data fusion while reducing dimensions and safeguarding privacy during the learning process. These processed data sets serve as inputs for cyber-attack detection, employing the quantum deep neural network method. Their solution underwent testing on two datasets, namely the WUSTL-EHMS [15] and ICU datasets [17], achieving an accuracy and detection rate exceeding 99%. Remarkably, their solution surpasses the outcomes of prior research efforts in this field.
@@
In addition, there have been proposals for IDS solutions that utilize a FL approach to strengthen privacy in the medical field. For instance, the authors of [35] suggest implementing an IDS to improve the security of the IoMT. The authors suggest implementing federated transfer learning, which utilizes DNN, in order to facilitate collaborative training of the cloud and edge models. The CICIDS2017 dataset is employed for the purpose of conducting experiments [24]. A multitude of metrics are employed in the performance evaluation process, encompassing accuracy, precision, detection rate, F1-score, training time, and testing time. The proposed model is evaluated in comparison to  centralized learning approaches, including Stochastic Gradient Descent (SGD), Deep Belief Network, and Support Vector Machine (SVM), in order to establish a benchmark. The obtained results indicate the effectiveness of the proposed model in terms of its ability to generalize and learn incrementally, outperforming the performance of other baseline ML/DL algorithms commonly used in traditional centralized learning approaches.

### v2

@@ Abstract
The Internet of Medical Things (IoMT) transcends traditional medical boundaries, enabling a transition from reactive treatment to proactive prevention. This innovative method revolutionizes healthcare by facilitating early disease detection and tailored care, particularly in chronic disease management, where IoMT automates treatments based on real-time health data collection. Nonetheless, its benefits are countered by significant security challenges that endanger the lives of its users due to the sensitivity and value of the processed data, thereby attracting malicious interests. Moreover, the utilization of wireless communication for data transmission exposes medical data to interception and tampering by cybercriminals. Additionally, anomalies may arise due to human error, network interference, or hardware malfunctions. In this context, anomaly detection based on Machine Learning (ML) is an interesting solution, but it comes up against obstacles in terms of explicability and privacy protection. To address these challenges, a new framework for Intrusion Detection Systems (IDS) is introduced, leveraging Artificial Neural Networks (ANN) for intrusion detection while utilizing Federated Learning (FL) for privacy preservation. Additionally, eXplainable Artificial Intelligence (XAI) methods are incorporated to enhance model explanation and interpretation. The efficacy of the proposed framework is evaluated and compared with centralized approaches using multiple datasets containing network and medical data, simulating various attack types impacting the confidentiality, integrity, and availability of medical and physiological data. The results obtained offer compelling evidence that the FL method performs comparably to the centralized method, demonstrating high performance. Additionally, it affords the dual advantage of safeguarding privacy and providing model explanation while adhering to ethical principles.
@@
Internet of Medical Things Intrusion Detection System Machine Learning Federated Learning eXplainable Artificial Intelligence Security Privacy.
@@ 1 INTRODUCTION
Internet of Things is a technology that revolutionizes the field of information science by incorporating sensors associated with objects to collect data. Its ability to obtain information anywhere and anytime has led to its integration into various sectors, including the healthcare domain known as IoMT. Equipped with sensors and actuators, medical devices facilitate continuous, remote, and real-time collection of physiological data, such as glucose levels, body temperature, and heart rate, allowing constant health monitoring. The enhancement in both the quantity and quality of the amassed data serves to optimize treatment efficacy, mitigate medical inaccuracies, and expedite early disease detection. This transformation means a transition from curative to preventive healthcare, which considerably increases the chances of patient recovery.
@@
Moreover, leveraging the gathered health data, automatic treatments can be administered to patients with chronic conditions through actuators. For example, diabetic patients can receive automated insulin injections based on their blood glucose levels. Similarly, individuals with irregular heart rhythms can be administered electrical shocks through pacemakers, while those with neurological disorders can benefit from simulated brain activity via Deep Brain Implants, enhancing the overall well-being and quality of life for patients with chronic illnesses. The health information collected can be stored on cloud servers or hospital databases for in-depth analysis, harnessing the power of AI-assisted healthcare under the supervision of healthcare professionals, often referred to as Healthcare 4.0.
@@
Notwithstanding the myriad benefits offered by IoMT, it faces significant security and privacy challenges, as evidenced by the alarming statistic that indicates a 77% increase in malware attacks on IoT devices in the first six months of 2022 [1]. The utilization of wireless communication for transmitting data exposes it to Man-In-The-Middle (MITM) attacks. Furthermore, other sources of anomalies may affect data integrity due to human errors during data processing, network interference during transmission, or malfunctions occurring at the medical equipment level [2]. These anomalies can compromise the integrity, confidentiality, and availability of crucial medical data. Such anomalies could lead to misdiagnoses and treatment errors, potentially resulting in tragic consequences.
@@
To mitigate these concerns regarding anomalies, ML has been proposed as a solution. Considering that IoMT systems generate large amounts of data, they can help ML models to distinguish between normal and abnormal behaviour. This capability facilitates the detection of anomalies and is effective against zero-day and new attacks. These detection systems can operate in real-time, and with advancements in Deep Learning (DL), the process of attribute selection and image processing [3] becomes automated.
Despite their potential benefits, ML-based security solutions face several challenges that must be addressed to ensure their effective, ethical, and regulatory-compliant deployment. A significant limitation lies in the integration of IDS based on ML into centralized architectures. While such architectures streamline data processing and model training, they raise critical concerns about data privacy and security. Sharing sensitive information across a centralized system may violate user privacy, particularly in sectors like healthcare, where patient data is highly confidential. Moreover, the central node itself represents a single point of failure: if compromised, it could jeopardize the entire system, leading to catastrophic consequences. Additionally, centralized architectures are prone to latency issues, as all data must be transmitted to and processed by the central node. This can hinder network scalability, limit computational capacity, and create bottlenecks, especially as data volumes grow.
Another major challenge is the inherent opacity of many ML models, often referred to as ’black-box’ models. These models lack transparency in their decision-making processes, making it difficult for stakeholders to understand how predictions or classifications are derived. This issue is particularly critical in regulated industries such as healthcare, where international standards like the Health Insurance Portability and Accountability Act (HIPAA) in the United States mandate explainability and transparency in outcomes. However, HIPAA is not the only standard to consider. International frameworks such as the General Data Protection Regulation (GDPR) in the European Union, ISO/IEC 27001 for information security management, and ISO/IEC 27701—an extension of ISO/IEC 27001 specifically focused on privacy information management—provide robust guidelines for data protection, privacy, and ethical AI deployment. The World Health Organization (WHO) has also emphasized the importance of transparency, explainability, and intelligibility in AI systems used in healthcare, highlighting that AI models must be interpretable by medical professionals, regulators, and patients to foster trust and ensure ethical deployment [4]. ISO/IEC 27701 complements ISO/IEC 27001 by adding specific requirements for privacy management, making it an essential tool for organizations seeking to align their security practices with international best practices in data privacy. Adopting these standards is crucial for organizations operating globally, as they ensure compliance with diverse regulatory requirements and foster user trust.
Furthermore, the ethical implications of ML-based security solutions are often overlooked in current discussions. Considerations such as fairness, accountability, and bias mitigation are critical to ensuring that these technologies do not inadvertently harm individuals or communities. For example, biased training data can lead to discriminatory outcomes, while a lack of accountability mechanisms can make it difficult to assign responsibility for errors or misuse. Addressing these ethical challenges requires a multidisciplinary approach, involving not only technical solutions but also input from ethicists, policymakers, and end-users. By integrating ethical principles into the design and implementation of ML systems, organizations can develop more trustworthy and socially responsible solutions.
@@
This article aims to introduce a framework to enhance the security of IoMT systems through the design of an IDS based on ML. The proposed solution involves the utilization of FL as a training methodology, allowing the sharing of locally trained model weights on end-devices instead of raw data. This approach preserves data privacy in alignment with regulations such as GDPR and HIPAA, while the distributed nature of FL mitigates the single point of failure associated with a centralized structure. By opting to share model weights rather than raw data, numerous key challenges are addressed. This reduces bandwidth consumption and alleviates network congestion, thereby facilitating system scalability. Adopting FL enhances data confidentiality, reduces the potential risks associated with a centralized model, and increases overall system efficiency and robustness.
From an ethical perspective, the proposed framework addresses critical concerns such as fairness, accountability, and bias mitigation. By leveraging FL, the framework ensures that sensitive data remains on local devices, reducing the risk of biased outcomes that can arise from centralized data collection. This decentralized approach allows for the inclusion of diverse datasets, promoting fairness and reducing the likelihood of discriminatory results. Furthermore, the integration of XAI enhances the transparency and interpretability of the ML models, enabling stakeholders—including patients, model designers, and regulators—to understand how decisions are made. This transparency fosters accountability, as it becomes easier to identify and address potential biases or errors in the system. By providing clear and understandable explanations of the model’s predictions, XAI ensures compliance with regulatory requirements that mandate explainability, such as those outlined in HIPAA, GDPR, WHO and ISO.
Through transparent explanations, a proven track record of reliable performance, and the provision of detection history in percentage form, this framework cultivates trust among users, facilitating confidence in the predictive capabilities of the system. By offering intuitive and accessible insights, even non-technical users can access results and track the model’s performance over time. The ethical design of the framework ensures that it not only meets technical and regulatory standards but also aligns with societal values, promoting the responsible and equitable use of AI in healthcare and beyond.
The contributions of the proposed solution can be summarized as follows:
@@
    The integration of XAI methods enhances transparency and interpretability, ensures regulatory compliance, and aids model designers in optimizing performance, while the demonstrated effectiveness of the detection history strengthens user confidence in the system.
@@
    The proposed solution’s performance is thoroughly evaluated and compared with the centralized method, demonstrating its effectiveness.
@@
    The evaluation of the solution proposed is conducted on four distinct datasets containing network and medical data, further validating its applicability and robustness.
  6. 6.
    The framework addresses ethical considerations such as fairness, accountability and bias mitigation through FL and XAI, ensuring privacy, transparency, and compliance with regulations.
The article is structured as follows: Section 2 presents the background and reviews related work. Section 3 details the methodology for the proposed IDS model. Section 4 describes the experimental setup and presents the results. Section 5 compares the proposed framework with prior studies. Section 6 discusses the contributions and implications of the work. Finally, Section 7 concludes the study and outlines future research directions.
@@ 2.1 Background
Shifting to ML, which can be categorized into supervised, unsupervised, and semi-supervised learning paradigms. Supervised ML involves training models on labeled data for classification or regression tasks, capturing relationships between inputs and output data [6]. Unsupervised learning explores patterns in unlabeled data [7], while semi-supervised learning leverages both labeled and unlabeled datasets to improve model performance and generalization [8]. The response time of an IDS spans the spectrum from real-time detection to a predefined time interval. Upon detecting an attack, the system initiates a response, which may be passive, involving the notification of the administrator regarding the presence of an attack, or active, entailing actions like communication blockage. The architecture can take different forms, such as Centralized, Fully-Distributed, and Partially Distributed IDS with hierarchical reporting mechanisms.
@@
Introducing FL, a distributed ML approach that preserves privacy by transmitting models instead of raw data. In FL, a global model is constructed by leveraging the participation of multiple devices over communication rounds. A subset of clients is chosen at the beginning of each round, updating the model independently using locally stored data. The server aggregates these models to create an enhanced global model, iterating until convergence. Two types of FL emerge: Cross-Silo FL, involving distributed data centers, and Cross-Device FL, encompassing mobile devices and IoT equipment [9]. FL encompasses various categorizations depending on the characteristics of the data distribution. These include Horizontal FL, where data across multiple devices share identical features but possess distinct samples; Vertical FL, where data across diverse entities exhibit differing features but may share common identifiers; and Federated Transfer Learning, which extends conventional Transfer Learning to the federated setting, allowing the transfer of knowledge across distributed data sources. [10], each contributing to data privacy and learning process improvement.
Table 1: Notation Table
Notation Description
$TP$ True Positives
$FP$ False Positives
$TN$ True Negatives
$FN$ False Negatives
$IoMT$ Internet of Medical Things
$FL$ Federated Learning
$XAI$ eXplainable Artificial Intelligence
$IDS$ Intrusion Detection System
$HIPAA$ Health Insurance Portability and Accountability Act
$WHO$ World Health Organization
$GDPR$ General Data Protection Regulation
$ISO$ International Organization for Standardization
$IEC$ International Electrotechnical Commission
$ML$ Machine Learning
$DL$ Deep Learning
$ANN$ Artificial Neural Network
$DNN$ Deep Neural Network
$SVM$ Support Vector Machine
$RF$ Random Forest
$KNN$ K-Nearest Neighbors
$CNN$ Convolutional Neural Network
$LSTM$ Long Short-Term Memory
$XSS$ Cross-Site Scripting
$TTL$ Time To Live
$U2R$ User to Root
$R2L$ Remote to Local
$ACCS$ Australian Centre for Cyber Security
$EHMS$ Enhanced Healthcare Monitoring System
$EMR$ Electronic Medical Record
$EHR$ Electronic Health Record
$CISO$ Chief Information Security Officer
$SHAP$ Shapley Additive Explanations
$MITM$ Man-In-The-Middle
$CSV$ Comma Separated Values
$AUC$ Area Under the Curve
$ROC$ Receiver Operating Characteristic
$LIME$ Local Interpretable Model-Agnostic Explanations
$GRU$ Gated Recurrent Unit
$PCA$ Principal Component Analysis
$AE$ Auto Encoder
$LRGU$ Logistic Redundancy Coefficient Upweighting
$MIFS$ Mutual Information Feature Selection
$GWO$ Grey Wolf Optimization
$SGD$ Stochastic Gradient Descent
$MCPS$ Medical Cyber-Physical Systems
$DoS$ Denial of Service
$DDoS$ Distributed Denial of Service
$HFL$ Hierarchical FL
$SMOTE$ Synthetic Minority Oversampling Technique
$DNS$ Domain Name System
$SGRU$ Sliced Gated Recurrent Unit
$sSAE$ Stacked Sparse Autoencoder
$ICU$ Intensive Care Unit
$TCP$ Transmission Control Protocol
$UDP$ User Datagram Protocol
$FTP$ File Transfer Protocol
$SSH$ Secure Shell
$HTTP$ Hypertext Transfer Protocol
$MLP$ Multi-Layer Perceptron
$LR$ Logistic Regression
$RFE$ Recursive Feature Elimination
$NB$ Naive Bayes
$PSO$ Particle Swarm Optimization
$NIDS$ Network-based IDS
$HIDS$ Host-based IDS
$DT$ Decision Tree
$SRU$ Simple Recurrent Unit
$Dintpkt$ Destination Inter Packet
$dstjitter$ Destination Jitter
$dstload$ Destination Load
$srcload$ Source Load
$dst\_port$ destination ports
$src\_port$ source port
@@
However, ML models are often considered black-box. To address this, Van Lent introduced XAI in 2004 [11]. XAI comprises methodologies aiding researchers in comprehending and gaining trust in ML model outcomes. A comprehensive AI model can be characterized with a three-dimensional description [12]. Explainability refers to the capacity to articulate the learning model’s processes, interpretability provides insight into the model’s operation, and transparency denotes inherent understandability without user intervention. In the domain of explainability, transparent models, like Decision Tree (DT), embody ante-hoc explainability, while opaque models, like Deep Neural Network (DNN), require post-hoc explainability. Nevertheless, there are scenarios where even transparent models with seemingly straightforward rules require post-hoc elucidation. Despite the complexity of opaque models, they often outperform transparent ones, leading to a trade-off between performance and explainability. Post-hoc methods operate on both local and global scop, employing techniques such as feature relevance analysis, which assesses the significance of each input feature in the model’s predictions. Surrogate models contribute to interpretability by simplifying the primary model through methods like probing with local changes or utilizing its structure. Additionally, representative examples, drawn from the model’s training set, showcase significant levels of confidence in their classification towards a particular class [13].
@@ 2.2 Related Work
Initially, centralized IDS solutions utilizing ML for anomaly detection are reviewed. Subsequently, an examination is conducted on solutions employing FL for intrusion detection. Finally, detection systems based on ML, incorporating XAI, are explored.
@@
The research presented in [16] introduced a methodology aimed at detecting anomalies and cyberattacks within healthcare systems. Their approach involves optimal feature selection using Recursive Feature Elimination (RFE), a process that iteratively eliminates the least relevant features using Logistic Regression (LR) or XGBoost Regressor methods. Subsequently, they employ a Multi-layer Perceptron (MLP) with parameter optimization for attack detection. Before applying their method, they conducted data preprocessing, replacing NaN values with column means and utilizing label encoding to transform categorical values into numerical ones. Performance evaluation was conducted using various metrics including accuracy, precision, recall, and f1-score. Their methods were applied to IoMT datasets such as WUSTL-EHMS [17], ECU-IoHT [18], Intensive Care Unit (ICU) [19], and ToN-IoT [20]. Furthermore, they compared their results with previous studies. The findings demonstrate that the combination of XGB Regressor for feature selection and optimized parameter settings for MLP yielded the most promising results among the tested methodologies.
In the work by [21] introduced an IDS designed specifically for connected healthcare systems. Their proposed methodology involves the utilization of Stacked Sparse Autoencoder (sSAE) to reduce dimensionality and the memory required for computing the covariance matrix. Subsequently, they employ the Sliced Gated Recurrent Unit (SGRU), a parallelized version of the RNN, achieved through processing segmentation. To measure the performance of their solution, the authors employed the AWID dataset [22]. However, recognizing the imbalance in the dataset, they solved this problem by applying the Synthetic Minority Oversampling Technique (SMOTE) to augment the original dataset with synthetic data. In their comparative analysis, the authors benchmarked their solution against DNN, Random Forest (RF), Long Short Term Memory (LSTM), LA-SMOTE-GRU [23], Auto Encoder (AE), XGBoost hybrid model, and LSTM hybrid model. Notably, their results demonstrate that their proposed solution outperforms these existing models, achieving superior performance metrics while simultaneously reducing model size and processing time.
@@
The authors of [29] have introduced an IDS tailored for the IoMT. Their innovative methodology incorporates PSO for feature selection followed by the utilization of DNN for intrusion detection within their proposed system. In their evaluation, the authors tested their solution on a dataset encompassing both network and medical features. To comprehensively assess their approach, they compared the results of their PSO-DNN solution with various ML and DL models, including LR, KNN, DT, RF, SVM, CNN, and LSTM on WUSTL-EHMS datasets [17]. For this comparative analysis, accuracy, precision, recall, and f1-score were utilized as evaluation metrics. Their findings revealed that the DL models surpassed the performance of traditional ML models. Furthermore, the PSO-DNN solution demonstrated a superior performance relative to conventional ML models and outperformed state-of-the-art approaches in intrusion detection within the IoMT domain.
@@
The IDS proposed in [33] is tailored for detecting data flow modifications within multi-cloud healthcare systems, comprising gateways, edge cloud, and a core cloud. Their proposed approach involves leveraging Deep Hierarchical Stacked Neural Networks. This methodology entails reusing layers trained at the edge cloud level and merging them at the core cloud level to create a pre-trained model. To validate their solution, they conducted testing on UNSW-BOT-IoT [34] and UNSW-NB15 [28] datasets, including one generated by the authors. A comparative analysis was performed between their model, which incorporates reusing the trained layers, and a model that does not. Their findings indicate that their solution enhances accuracy while reducing training time.
@@
Finally, the comparative study by [35] emphasizes the impact of combining network and medical attributes for intrusion detection in IoMT. Initially, they undertook the creation and collection of both medical and network data, along with simulating a MITM attack. Subsequently, they tested various ML algorithms, including ANN, RF, KNN, and SVM, based on metrics such as AUC, accuracy, and execution time. The results indicate that the performance of ML algorithms utilizing the combination of network and medical attributes yields superior results compared to using them separately.
@@
Continuing our exploration, attention now turns towards IDS solutions harnessing the power of FL. A noteworthy example is found in the work of [36], where the authors present a framework for storing and privacy protection of data used to detect cyber-attacks within IoMT systems. Their approach involves distributing heterogeneous data across cloud nodes with privacy protection measures. They further suggested merging data from diverse sources using the differential privacy contractive deep autoencoder. This technique facilitates data fusion while reducing dimensions and safeguarding privacy during the learning process. These processed data sets serve as inputs for cyber-attack detection, employing the quantum DNN method. Their solution underwent testing on two datasets, namely the WUSTL-EHMS [17] and ICU datasets [19], achieving an accuracy and detection rate exceeding 99%. Remarkably, their solution surpasses the outcomes of prior research efforts in this field.
@@
In addition, there have been proposals for IDS solutions that utilize a FL approach to strengthen privacy in the medical field. For instance, the authors of [37] suggest implementing an IDS to improve the security of the IoMT. The authors suggest implementing federated transfer learning, which utilizes DNN, in order to facilitate collaborative training of the cloud and edge models. The CICIDS2017 dataset is employed for the purpose of conducting experiments [26]. A multitude of metrics are employed in the performance evaluation process, encompassing accuracy, precision, detection rate, F1-score, training time, and testing time. The proposed model is evaluated in comparison to centralized learning approaches, including Stochastic Gradient Descent (SGD), Deep Belief Network, and SVM, in order to establish a benchmark. The obtained results indicate the effectiveness of the proposed model in terms of its ability to generalize and learn incrementally, outperforming the performance of other baseline ML/DL algorithms commonly used in traditional centralized learning approaches.